OpenStudy (anonymous):

Brooke wants to have equal amounts of room on all four sides of her pool table. A 4ft by 7ft table is placed in the center of a rectangular room that has an area of 238 square feet. What is the width of the clear space Brooke will have on all sides of her table? Show your work.

OpenStudy (anonymous):

5 feet x=clear space solve the following for x (2 x + 7) (2 x + 4) = 238

hero (hero):

Area of clear space = 238 - 4*7 = 238 - 28 = 210

OpenStudy (across):

The room is rectangular.

hero (hero):

across, area = xy

OpenStudy (across):

robtobey is using the same variable for both the length and the height.

hero (hero):

Yeah, that can be a problem: (2x+4)(2y+7) = 238

hero (hero):

Oh...because x and y are apparently the same on all four sides

hero (hero):

so x = y

hero (hero):

Maybe robtobey was on the right track after all

OpenStudy (anonymous):

I hope. Recheck the first problem sentence.

hero (hero):

Yeah, good job rob

OpenStudy (anonymous):

Thank you.